# Phi-3 Technical Report: A Highly Capable Language Model Locally on Your Phone

### v2

@@ Appendix A Example prompt for benchmarks
Appendix B Authors
Marah Abdin Russell J. Hewett Olatunji Ruwase
Sam Ade Jacobs Jamie Huynh Olli Saarikivi
Ammar Ahmad Awan Mojan Javaheripi Amin Saied
Jyoti Aneja Xin Jin Adil Salim
Ahmed Awadallah Piero Kauffmann Michael Santacroce
Hany Awadalla Nikos Karampatziakis Shital Shah
Nguyen Bach Dongwoo Kim Ning Shang
Amit Bahree Mahmoud Khademi Hiteshi Sharma
Arash Bakhtiari Lev Kurilenko Xia Song
Harkirat Behl James R. Lee Masahiro Tanaka
Alon Benhaim Yin Tat Lee Xin Wang
Misha Bilenko Yuanzhi Li Rachel Ward
Johan Bjorck Chen Liang Guanhua Wang
Sébastien Bubeck Weishung Liu Philipp Witte
Martin Cai Eric Lin Michael Wyatt
Caio César Teodoro Mendes Zeqi Lin Jiahang Xu
Weizhu Chen Piyush Madan Can Xu
Vishrav Chaudhary Arindam Mitra Sonali Yadav
Parul Chopra Hardik Modi Fan Yang
@@ Appendix B Authors
Gustavo de Rosa Anh Nguyen Donghan Yu
Matthew Dixon Barun Patra Chengruidong Zhang
Ronen Eldan Daniel Perez-Becker Cyril Zhang
Dan Iter Heyang Qin Jianwen Zhang
Amit Garg Thomas Portet Li Lyna Zhang
Abhishek Goswami Reid Pryzant Yi Zhang
Suriya Gunasekar Sambuddha Roy Yue Zhang
Emman Haider Marko Radmilac Yunan Zhang
Junheng Hao Corby Rosset Xiren Zhou

### v3

@@ Appendix A Example prompt for benchmarks
Appendix B Authors (alphabetical)
Marah Abdin Jamie Huynh Olli Saarikivi
Sam Ade Jacobs Dan Iter Amin Saied
Ammar Ahmad Awan Russell J. Hewett Adil Salim
Jyoti Aneja Mojan Javaheripi Michael Santacroce
Hany Awadalla Xin Jin Shital Shah
Ahmed Awadallah Nikos Karampatziakis Ning Shang
Nguyen Bach Piero Kauffmann Hiteshi Sharma
Amit Bahree Mahoud Khademi Swadheen Shukla
Arash Bakhtiari Dongwoo Kim Xia Song
Jianmin Bao Lev Kurilenko Yin Tat Lee
Harkirat Behl Yuanzhi Li Andrea Tupini
Alon Benhaim Yunsheng Li Xin Wang
Misha Bilenko Chen Liang Guanhua Wang
Johan Bjorck Lars Liden Lijuan Wang
Sébastien Bubeck Eric Lin Chunyu Wang
Caio César Teodoro Mendes Zeqi Lin Yu Wang
Martin Cai Weishung Liu Rachel Ward
Qin Cai Mengchen Liu Wen Wen
Vishrav Chaudhary Ce Liu Philipp Witte
Weizhu Chen Chong Luo Haiping Wu
Yi-Ling Chen Li Lyna Zhang Michael Wyatt
Yen-Chun Chen Piyush Madan Bin Xiao
Dongdong Chen David Majercak Jiahang Xu
Dong Chen Matt Mazzola Can Xu
Parul Chopra Arindam Mitra Weijian Xu
Xiyang Dai Hardik Modi Sonali Yadav
Gustavo de Rosa Anh Nguyen Fan Yang
@@ Appendix B Authors (alphabetical)
Matthew Dixon Barun Patra Jianwei Yang
Ronen Eldan Daniel Perez-Becker Yifan Yang
Victor Fragoso Thomas Portet Donghan Yu
Mei Gao Reid Pryzant Lu Yuan
Min Gao Heyang Qin Chengruidong Zhang
Jianfeng Gao James R. Lee Cyril Zhang
Abhishek Goswami Marko Radmilac Jianwen Zhang
Suriya Gunasekar Corby Rosset Yi Zhang
Emman Haider Sambudha Roy Yunan Zhang
Junheng Hao Olatunji Ruwase Xiren Zhou
Appendix C Acknowledgements
We would like to thank Zhuohan Li, Simon Mo, and Kaichao You from UC Berkeley for sharing their insights on the vLLM kernel.